# Building Bridges, Not Walls: Advancing Interpretability by Unifying Feature, Data, and Model Component Attribution

### v2

@@ 5 Understanding Component Attributions
Component attribution, an emerging approach within mechanistic interpretability, seeks to understand AI models by reverse engineering their internal mechanisms into interpretable algorithms. Operating primarily at test time for model inference, it quantifies how each model component contributes to a model output through an attribution score . Components can be defined flexibly across different scales - from individual neurons and attention heads to entire layers and circuits (subnetworks). By identifying components responsible for specific behaviors, this approach enables deeper model understanding and targeted model editing. Like feature and data attribution, component attribution methods fall into three categories: perturbation-based, gradient-based, and linear approximation approaches. Below we examine key methods from each category, with additional details provided in Appendix E.
5.1 Perturbation-Based Methods
In component attribution, perturbation-based methods are fundamentally quite similar to perturbation-based methods in feature and data attribution. Components of the model, whether neurons, circuits, or layers, are similarly perturbed to measure their effect on model behavior. Generally, the perturbations are chosen carefully to attempt to localize behaviors related to specific tasks or concepts.
@@ 5.1 Perturbation-Based Methods
Causal Mediation Analysis (Pearl, 2022; Vig et al., 2020) is based on the abstraction of models to causal graphs. These graphs consist of nodes, which can be components such as neurons, circuits, attention heads, or layers, and directed edges that represent the causal relationships between nodes. Causal mediation analysis is defined by an input cause and an output effect that is mediated by intermediate causal nodes between and . By perturbing these intermediate components, , changes in can be measured to get attribution scores . These indirect effects are often measured counterfactually in order to calculate each component’s contribution towards a particular behavior, such as a correct factual prediction. To do so, the activations of all intermediate components are measured during three separate runs: a clean run with no perturbations, a corrupted run where intermediate activations are perturbed, and a corrupted-with-restoration run that measures whether a single component can restore the prediction. The corrupted run can be repeated multiple times with different random noise added to obtain a more robust attribution score. This analysis is frequently referred to as causal tracing (Meng et al., 2022) or activation patching, and also path patching (Wang et al., 2022) when patching is applied to paths connecting components. By comparing the outputs of the clean and corrupted run, or by looking at the corrupted-with-restoration run, one is able to find the specific mediator components that are either sufficient or necessary for the desired behavior. By changing the dataset, metric, and causal mediators, we can model the relationship between each component and various model behaviors.
Game-Theoretic Component Attribution (Ghorbani and Zou, 2020) follows a similar approach to game-theoretic methods in feature and data attribution to quantify the contributions of each neuron to the model’s performance. These methods take into account the interactions between neurons by modeling neurons as players in a cooperative game to fairly distribute contributions. In particular, Neuron Shapley (Ghorbani and Zou, 2020) extends prior works on Shapley values to component attribution, ensuring computational feasibility through sampling-based approximations and a multi-armed bandit algorithm that efficiently identifies neurons with large attribution scores.
Mask Learning and Subnetwork Probing (Csordás et al., 2020; Cao et al., 2021) adopts a similar concept to feature attribution, attempting to approximate either the model’s or a probe’s performance on a given task by searching for a subnetwork or components that equivalently perform that task. More specifically, subnetwork probing optimizes a mask for the weights of the model, essentially pruning the model, by performing gradient descent on a continuous relaxation of searching for the subnetwork of a model that performs the task of interest. Thus, behavior can be attributed to the parts of the network that are not masked out.
5.2 Gradient-Based Component Attribution
@@
To further decrease the computational complexity of component attribution methods, researchers have developed alterations of causal tracing that leverage gradient-based approximations requiring only two forward passes and a single backward pass to generate attributions. Attribution patching (Nanda, 2023) is the simplest gradient-based approximation of causal tracing. Intuitively, attribution patching leverages a linear approximation of the model for the corrupted prompts and measures the local change when patching a single activation from the corrupted to clean input. This is achieved by computing the backward pass for the corrupted output with respect to the patching metric and storing the gradients with respect to the activations. Note that for feature and data attributions, gradients are taken with respect to the input features or training data, not the model activations. Finally, the method takes the difference between the clean and corrupted activations and multiplies it by the cached gradients to obtain attribution scores.
5.3 Linear Approximation for Component Attribution
@@
Given the rapid increase in model size and the combinatorial nature of searching for effective components, component attribution also employs linear approximations like LIME and Datamodels. COAR (Shah et al., 2024) attempts to decompose model behavior in terms of various model components by predicting the counterfactual impact of ablating each component, similar to many forms of causal mediation analysis. Given the computational complexity of this problem, they employ linear approximations by assigning scores to each component of a model and estimating the counterfactual effect of removing sets of components by simply summing their corresponding scores. Thus, the complexity of relationships between components is abstracted away through the linear approximation.
6 Position and Contributions
Feature, data, and component attribution methods have largely been studied as separate problems, resulting in the parallel development of similar methods from different communities with distinct terminologies. We argue that these methods can be unified into a holistic view. Having demonstrated their methodological similarities across three types of attribution, we now summarize their common concepts and challenges and identify promising research directions through cross-aspect knowledge transfer. We believe that our unified view will bridge the current fragmented landscape, make the field more accessible to newcomers, and help advance research in interpretability and beyond.
6.1 Common Concepts of Attribution Methods
@@
As we discussed in the previous sections, attribution methods across features, data, and components can be categorized into: perturbation-based, gradient-based, and linear approximation. We provide a detailed discussion of these categories in Appendix B with Table 1 summarizing all the methods we discussed. Beyond algorithmic similarities, conceptual ideas also transfer across aspects of attribution. One example is the deliberate introduction of randomness and smoothing to enhance attribution robustness. This idea has proven effective across three attribution types through random noisy samples used by SmoothGrad for feature attribution, attribution results ensembled over multiple retrainings by TRAK for data attribution, and aggregated results from multiple corrupted runs in causal mediation analysis for component attribution. Another example is tracking and aggregating results along paths, as Integrated Gradients along paths from base input to target input in feature attribution, TracIn tracing training paths to reveal dynamic data influences in data attribution, and path patching tracking component effects along residual stream paths in component attribution. These shared concepts highlight the fundamental connections of attribution methods.
6.2 Common Challenges of Attribution Methods
Attribution methods also face common challenges that impact their reliability and practical utility, which we briefly discuss below and extend in Appendix F.
@@ 6.2 Common Challenges of Attribution Methods
Computational Challenges present substantial barriers preventing attribution methods from being applied to large models. These challenges appear in all three types of attributions, rooted in their shared technical methods. For perturbation-based methods, the curse of dimensionality makes it intractable to comprehensively analyze high-dimensional inputs, large training datasets, and models with numerous components. Gradient-based methods offer more practical computational costs, except when sophisticated gradient computations are needed, such as aggregating mang gradients or computing second-order Hessian matrices. Linear approximation methods also face challenges when numerous data points and model evaluations are required to establish sufficient data for learning accurate linear models.
Consistency Challenges refer to the variability in attribution results across multiple runs of the same method with different random seeds, making it challenging to establish stable interpretations and evaluations. This challenge is prevalent across all three types of attribution due to multiple sources of randomness in their common techniques, including sampling, learning processes with stochastic optimization, and also non-trivial hyperparameters. While some gradient-based methods can produce consistent results in a single run when they do not involve sampling or approximations for computationally intensive operations, the consistency between different gradient-based methods varies considerably, which is a problem for all three method categories and leads to the following evaluation challenges.
@@
Evaluation Challenges arise in all three types of attribution due to multiple factors. Inconsistent results make it difficult to reliably compare methods and determine their relative accuracy. This challenge is particularly evident in gradient-based feature attribution methods. As demonstrated by Adebayo et al. (2018), they can produce contradictory attribution results and sometimes perform no better than random baselines, making them difficult to evaluate fairly. For existing evaluation metrics, counterfactual evaluation could provide more rigorous validation, but computational constraints often make this approach impractical. Task-specific evaluations offer easier alternatives but frequently lack generalizability across different contexts. Human evaluation, despite being considered the gold standard, faces scalability issues and potential biases. The diversity of evaluation metrics and their varying definitions of importance make the evaluation more challenging. An attribution result may perform well by one metric but poorly according to another. These challenges emphasize the pressing need for developing more reliable and practical attribution evaluation metrics.
6.3 Cross-Aspect Attribution Innovation
@@
The connections among feature, data, and component attributions discussed in the sections above suggest multiple promising directions for future research. One research direction is to leverage insights from one type of attribution to develop methods for another. This can be directly identified as filling in the empty cells in Table 1. For example, while the Shapley value has been successfully applied across all three types of attribution, many other game-theoretic notions have only been used in feature attribution and not for data and component attribution. In addition, some advanced gradient techniques are common for feature and data attribution methods, but not for component attribution. The Hessian matrix, for example, has been used to obtain second-order information in Integrated Hessians for feature attribution and extensively in all IF-related methods for data attribution, and can be explored for component attribution.
@@
Moreover, seeing the theoretical connections among feature, data, and component attributions enables us to draw inspiration from one area to advance our understanding of another as a whole. For example, we demonstrated that diverse feature attribution methods all perform local function approximation (§ 3.4). This framework can potentially also apply to data and component attributions. We know that feature attributions perform function approximation of the blackbox model’s predictions over the space of input features. One may hypothesize that data attributions perform function approximation of the model’s weights over the space of training data points and that component attributions perform function approximation of the model’s predictions over the space of model components. If so, function approximation may unify data and component attribution methods as well. Such theoretical unification may provide plentiful benefits to data and component attribution, including conceptual coherence, elucidation of method properties, theoretical simplicity, and clearer practical recommendations.
@@
Another research direction is to move towards more holistic analyses of model behavior. These attribution methods provide insight into model behavior through different lenses: input features, training data, and model components. Each type of attribution provides different and complementary information about model behavior. For example, for a given model prediction, feature attributions may not suggest that the model is relying on sensitive features to make predictions, but model component attributions may uncover a set of neurons that encode biased patterns. In this sense, focusing only on one type of attribution, i.e., studying only one part of the model, is insufficient to understand model behavior. Thus, future research may develop approaches to enable more comprehensive model understanding, such as understanding how to use different types of attribution methods together, the settings under which different attribution types may support or contradict one another, and the interactions between the three model parts (e.g., how patterns in the training data are encoded in model neurons).
6.4 Connections to Other Areas of AI
@@
Attribution methods also hold immense potential to benefit other AI areas. Especially with a unified view integrating feature, data, and component attribution, researchers can not only gain deeper insights of model behavior but also edit and steer models towards desired goals and improve model compliance with regulatory standards.
@@
Model Editing (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022, inter alia) focuses on precisely modifying models without retraining. It enables researchers to correct model mistakes, analogous to fixing bugs in software. This approach is particularly valuable for large language models (LLMs), which encode vast information in their parameters and are prohibitively expensive to retrain. It can be viewed as a downstream task of attribution methods. Once attribution methods locate an issue, editing methods can be applied to the problematic parts. While editing aligns most closely with component attribution, other attribution types serve essential complementary functions. Feature attribution identifies spurious correlations requiring correction, and data attribution reveals problematic training samples that influence model behavior. The unified attribution framework provides a holistic perspective that enables more efficient and accurate editing, especially when component attribution alone proves insufficient (Hase et al., 2024).
@@
Model Steering (Zou et al., 2023, inter alia) differs from model editing by integrating a steering vector into the model’s inference process rather than modifying model parameters. While editing focuses on specific knowledge modifications, steering guides model behavior at a higher level, such as enhancing truthfulness and harmlessness in LLMs. Similar to model editing, a unified attribution framework can significantly enhance steering by better localizing target components to steer and generating more effective steering vectors through relevant features and training data.
@@
Model Regulation (Oesterling et al., 2024, inter alia) is an emerging field examining the relationship between AI systems, policy, and societal outcomes. Regulation and policy frequently stress the need for transparency of AI systems as well as users’ right to an explanation. Attribution methods provide an avenue for practitioners to ensure that AI systems meet these legal and ethical requirements, by providing information about the overall AI system as well as specific input-output behavior. Feature attribution reveals input processing patterns, data attribution exposes training data influences, and component attribution illuminates architectural roles. This multi-faceted understanding enables more targeted and effective regulation. For example, when addressing biased behavior, feature attribution can be used to identify discriminatory input patterns, data attribution to trace problematic training samples or copyright infringements, and component attribution to locate architectural elements needing adjustment. These complementary perspectives provide the comprehensive understanding needed to guide model regulation toward desired societal outcomes.

### v3

@@ 3.3 Understanding component attributions (CA)
CA studies the role of internal model components in model outputs. It plays a central role in the emerging mechanistic interpretability research which seeks to understand AI models by reverse engineering their internal mechanisms into interpretable algorithms. CA operates primarily at test time for model inference, quantifying how each model component contributes to the output through an attribution score . Components can be defined flexibly - from individual neurons and attention heads to entire layers and circuits (subnetworks). Like FA and DA, CA leverages the same three techniques: perturbations, gradients, and linear approximation. We examine key CA methods below and provide additional details in Appendix E.
@@
Perturbation-based CA. Perturbation-based methods for CA are fundamentally similar to those for FA and DA. Components of the model, whether neurons, layers, or circuits, are perturbed to measure their effect on model behavior. Generally, the perturbations are chosen carefully to attempt to localize behaviors related to specific tasks or concepts. Causal mediation analysis (Pearl, 2022; Vig et al., 2020) is a perturbation-based method that leverages the abstraction of models to causal graphs, with graphs nodes corresponding to components, and graph edges represent the causal relationships between nodes. Causal mediation analysis is defined by an input cause and an output effect that is mediated by intermediate causal nodes, , between and . By perturbing , changes in can be measured to get attribution scores . Specifically, the effects of perturbations on component activations are measured during three separate runs: a clean run with no perturbations, a corrupted run where components are perturbed (may also be perturbing inputs to induce corrupted component activations), and a corrupted-with-restoration run that measures how much restoring a single component activation of the corrupted run by its original clean-run value can restore the output. The corrupted run can be repeated multiple times with different random noise added to obtain more robust CA scores. This analysis is frequently referred to as causal tracing (Meng et al., 2022) or activation patching, and path patching (Wang et al., 2022) when patching is applied to paths connecting components. Game-theoretic CA (Ghorbani and Zou, 2020) follows a similar approach to game-theoretic FA and DA. In particular, Neuron Shapley (Ghorbani and Zou, 2020) extends prior works on Shapley values to CA and mitigates computational cost through sampling and a multi-armed bandit algorithm to identifies neurons with large attribution scores. Mask Learning and Subnetwork Probing (Csordás et al., 2020; Cao et al., 2021) adopts a similar concept to FA mask learning, attempting to approximate either the model’s or a probe’s performance on a given task by searching for components that equivalently perform that task. More specifically, subnetwork probing optimizes a mask for the model weights as a continuous relaxation for searching, which essentially prunes the model.
@@
Gradient-based CA. To improve the computational efficiency of CA methods, alterations of causal tracing leverage gradient-based approximations to do only two forward passes and a single backward pass to generate attributions. Attribution patching (Nanda, 2023) approximates the local change when patching a single activation from the corrupted to clean input. This is achieved by computing the backward pass for the corrupted output with respect to the patching metric and storing the gradients with respect to the activations. Finally, the difference between the clean and corrupted activations is taken and multiplied by the cached gradients to obtain attribution scores.
@@
Linear approximation for CA. CA also employs linear approximations like LIME in FA and Datamodels in DA. COAR (Shah et al., 2024) attempts to decompose model behavior in terms of various components by predicting the counterfactual impact of ablating each component. Given the computational complexity, COAR employs linear approximations by assigning scores to each component and estimating the counterfactual effect of removing sets of components by summing their scores. Thus, the complexity of relationships between components is abstracted away.
3.4 Common concepts among attribution methods
@@
As demonstrated in the previous sections, FA, DA, and CA methods leverage three shared techniques: perturbations, gradients, and linear approximations. Additional details are discussed in Appendix B and a summary of the methods is provided in Table 1. Beyond shared techniques, conceptual ideas also transfer across the three types of attributions. One such idea is the deliberate introduction of randomness and smoothing to enhance attribution robustness. This idea has proven effective across all three attribution types, with SmoothGrad using random noisy samples to denoise gradients for FA, TRAK ensembling attributions over multiple retrainings for DA, and causal mediation analysis aggregating results from multiple corrupted runs for CA. Another shared conceptual idea is tracking and aggregating results along paths. For example, Integrated Gradients calculates the path integral from base input to target input in FA, TracIn traces training paths to reveal dynamic data influences in DA, and path patching tracks component effects along residual stream paths in CA. These shared concepts are high-level ideas, not specific algorithms, and the definitions of paths can be different, but they highlight fundamental connections among different types of attributions. For example, both Integrated Gradients and TracIn admit inspiration from the Fundamental Theorem of Calculus.
Common approaches have also been utilized to evaluate the quality of all three types of attribution (Agarwal et al., 2022; Hammoudeh and Lowd, 2024; Rai et al., 2024). Counterfactual evaluation is a widely used approach, which compares the attribution scores with the actual impact of removing or modifying inputs. For example, evaluate by measuring the model’s performance drop when the inputs with high attribution scores are removed. Task-specific evaluation is another common approach to measure the attribution quality by running specific downstream tasks with the attribution results. In this type of evaluation, one can use FA to help identify feature changes that can flip model outputs, apply DA to detect mislabeled training examples, and run CA to identify important components that allows for model pruning. Moreover, human evaluation in many cases can be a gold standard by relying on domain experts or users to assess the quality and interpretability of attributions. More details on these evaluation approaches and their pros and cons are provided in Section F.3.
4 Leveraging the unified perspective to advance research
Feature, data, and component attribution have been studied largely as separate research areas, resulting in the parallel development of similar methods from different communities with distinct terminologies. Having demonstrated the unified view of the three types of attribution, we now discuss how this unified view can help advance both interpretability and broader AI research. We point out open challenges as well as potential solutions as future research directions.
4.1 Advancing research in interpretability
4.1.1 Overcoming shared challenges among all attribution types
Attribution methods face several common challenges that need to be addressed to enhance their practicality and reliability. Our unified perspective provides clearer insights into the underlying causes of these challenges and their potential solutions. We discuss these key challenges and propose potential solutions both below and in greater detail in appendix F.
@@ 4.1.1 Overcoming shared challenges among all attribution types
Improving computational efficiency. Computational efficiency is a significant challenge for all three types of attribution methods, preventing them from being applied to large models (Adolfi et al., 2024; Bassan et al., 2024; Barceló et al., 2020). Our unified view reveals that this challenge is rooted in their shared techniques. For perturbation-based methods, the curse of dimensionality makes it intractable to comprehensively analyze high-dimensional inputs, large training datasets, and models with numerous components. Gradient-based methods offer more practical computational costs, except when sophisticated gradient computations are needed, such as aggregating gradients or computing second-order Hessian matrices. Linear approximation methods also face computational challenges when numerous data points and model evaluations are required to establish sufficient data for learning accurate linear models. Therefore, improving computational efficiency calls for more efficient sampling methods (e.g., Maximum Sample Reuse (Wang and Jia, 2023)) to avoid perturbing the entire input space in perturbation-based methods and to learn more accurate linear models for linear approximation methods. Also needed are more efficient gradient computations, especially for higher-order gradients like Hessians. Many works have sought to improve these aspects, with promising examples like Grosse et al. (2023) showing that gradient-based methods can be scaled to large language models (LLMs) with up to 52 billion parameters. However, there is still much room for improvement to make attribution methods more efficient. Also, a unified perspective may be developed for the corresponding computational problems (Adolfi et al., 2024; Bassan et al., 2024; Barceló et al., 2020).
Improving attribution consistency. Consistency challenges refer to the variability in attribution results across multiple runs of the same method with different random seeds, making it challenging to establish stable interpretations and evaluations (Jeyakumar et al., 2020). Our unified view reveals that this challenge is prevalent across all three attribution types due to multiple of randomness rooted in their common techniques, including sampling, learning processes with stochastic optimization, and non-trivial hyperparameters. We see the potential improvement for this challenge including producing more deterministic attribution methods that are less sensitive to random hyperparameters, provide provable guarantees of attribution stability or error bounds. While past works have made some progress in this direction, there were strong assumptions added to the setting, e.g., convexity (Koh and Liang, 2017). Thus, more efforts are needed to improve consistency in more realistic settings.
@@
Towards fairer and more practical evaluation. Evaluation challenges arise in all three types of attribution due to multiple factors. Inconsistent results make it difficult to reliably compare methods and determine their relative accuracy. This challenge is particularly evident in gradient-based FA methods. As demonstrated by Adebayo et al. (2018), they can produce contradictory attribution results and sometimes perform no better than random baselines, making them difficult to evaluate fairly. For the shared evaluation metrics among all three types of attribution discussed in Section 3.4, counterfactual evaluation could provide more rigorous validation, but computational constraints often make this approach impractical. Task-specific evaluations offer easier alternatives but frequently lack generalizability across different contexts. Human evaluation, despite being considered the gold standard, faces scalability issues and potential biases. The diversity of evaluation metrics and their varying definitions of importance make the evaluation more challenging. An attribution result may perform well by one metric but poorly according to another. These challenges emphasize the pressing need for developing more reliable and generalizable attribution evaluation metrics.
Altogether, these are open challenges that call for future research to develop more efficient and reliable attribution methods as well as rigorous and principled frameworks to evaluate them. Given the shared techniques and common challenges among different attribution types, ideas to address these challenges in one attribution type will likely also be relevant to others.
4.1.2 Cross-attribution innovation
@@
The connections among FA, DA, and CA discussed in the sections above naturally suggest promising research directions of leveraging insights from one type of attribution to develop methods for another. This can be directly identified as filling in the empty cells in Table 1. For example, while the Shapley value has been successfully applied across all three types of attribution, many other game-theoretic notions have only been used in FA and not for DA and CA yet. In addition, some advanced gradient techniques are common for FA and DA methods, but not for CA methods. The Hessian matrix, for example, has been used to obtain second-order information in Integrated Hessians for FA and extensively in all IF-related methods for DA, and can be explored for CA.
@@
Moreover, seeing the theoretical connections among FA, DA, and CA enables us to draw inspiration from one area to advance our understanding of another as a whole. For example, Section § 3.1.1 discusses how diverse FA methods all perform LFA. This framework may also apply to DA and CA. We know that FA performs function approximation of the blackbox model’s predictions over the space of input features. One may hypothesize that DAs perform function approximation of the model’s weights over the space of training data points and that CAs perform function approximation of the model’s predictions over the space of model components. If so, function approximation may unify DA methods and unify CA methods as well. As in the case of unifying FA methods, such theoretical unification may also provide many benefits to DA and CA, including conceptual coherence, elucidation of method properties, simplicity for theoretical analyses, and principled practical recommendations.
@@
Another research direction is to move towards more holistic analyses of model behavior. These attribution methods provide insight into model behavior through different lenses: input features, training data, and model components. Each type of attribution provides different and complementary information about model behavior. For example, for a given model prediction, FA may not suggest that the model is relying on sensitive features to make predictions, but CA may uncover a set of neurons that encode biased patterns. In this sense, focusing only on one type of attribution, i.e., studying only one part of the model, is insufficient to understand model behavior. Thus, future research may develop approaches to enable more comprehensive model understanding, such as understanding how to use different types of attribution methods together, the settings under which different attribution types may support or contradict one another, and the interactions between the three model parts (e.g., how patterns in the training data are encoded in model neurons).
4.2 Advancing research in broader AI
@@
Attribution methods also hold immense potential to benefit broader AI research for other applications. Especially with a unified view integrating FA, DA, and CA, researchers cannot only gain deeper insights of model behavior but also edit and steer models towards desired goals and improve model compliance with regulatory standards.
@@
Model editing (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022, inter alia) focuses on precisely modifying models without retraining. It enables researchers to correct model mistakes, analogous to fixing bugs in software. This approach is particularly valuable for LLMs, which encode vast information in their parameters and are prohibitively expensive to retrain. It can be viewed as a downstream task of attribution methods. Once attribution methods locate an issue, editing methods can be applied to the problematic parts. While editing aligns most closely with CA, other attribution types serve essential complementary functions. FA identifies spurious correlations requiring correction, and DA helps select proper data points to induce efficient editing. The unified attribution framework provides a holistic perspective that enables more efficient and accurate editing, especially when CA alone proves insufficient (Hase et al., 2024).
@@
Model steering (Zou et al., 2023; Marks and Tegmark, 2023; Bürger et al., 2024, inter alia) differs from model editing by integrating a steering vector into the model’s inference process rather than modifying model parameters. While editing focuses on specific knowledge modifications, steering guides model behavior at a higher level, such as enhancing truthfulness and harmlessness in LLMs. Similar to model editing, a unified attribution framework can significantly enhance steering by better localizing target components to steer and generating more effective steering vectors through relevant features and training data. Recently, steering methods that rely on more than one type of attribution methods have emerged Marks and Tegmark (2023); Bürger et al. (2024). In Appendix G, we provide a more detailed discussion of how a unified attribution framework can enhance both model editing and steering and provide a illustrative example of steering for truthfulness of LLMs.
@@
Model regulation (Oesterling et al., 2024, inter alia) is an emerging field examining the relationship between AI systems, policy, and societal outcomes. Regulation and policy frequently stress the need for transparency of AI systems as well as users’ right to an explanation. Attribution methods provide an avenue for practitioners to ensure that AI systems meet these legal and ethical requirements, by providing information about the overall AI system as well as specific input-output behavior. FA reveals input processing patterns, DA exposes training data influences, and CA illuminates architectural roles. This multi-faceted understanding enables more targeted and effective regulation. For example, when addressing biased behavior, FA can be used to identify discriminatory input patterns, DA to trace problematic training samples or copyright infringements, and CA to locate architectural elements needing adjustment. These complementary perspectives provide the comprehensive understanding needed to guide model regulation toward desired societal outcomes.